40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon May 4 10:38:52 2009, steve, Update, MOPA, laser power is dropped dtecup.jpg
    Reply  Mon May 4 16:49:56 2009, Alberto, Update, MOPA, laser power is dropped 
    Reply  Tue May 5 10:42:18 2009, steve, Update, MOPA, laser power is back uncloged.jpg
Message ID: 1547     Entry time: Tue May 5 10:42:18 2009     In reply to: 1542
Author: steve 
Type: Update 
Category: MOPA 
Subject: laser power is back 

Quote:

As PSL-126MOPA_DTEC went up, the power out put went down yesterday

 The NPRO cooling water was clogged at the needle valve. The heat sink temp was around ~37C

The flow-regulator  needle valve position is locked with a nut and it is frozen. It is not adjustable. However Jeenne's tapping and pushing down on the plastic hardware cleared the way for the water flow.

We have to remember to replace this needle valve when the new NPRO will be swapped in. I checked on the heat sink temp this morning. It is ~18C

There is condensation on the south end of the NPRO body, I wish that the DTEC value would just a little higher like 0.5V

The wavelenght of the diode is temp dependent: 0.3 nm/C. The fine tuning of this diode is done by thermo-electric cooler ( TEC )

To keep the diode precisely tuned to the absorption of the laser gain material the diode temp is held constant using electronic feedback control.

This value is zero now.

 

Attachment 1: uncloged.jpg  66 kB  | Hide | Hide all
uncloged.jpg
ELOG V3.1.3-